OpenStudy (anonymous):

A rectangular room must have a maximum area of 84 square feet. If the room is to be 8 feet wide, what is the longest the room can be?

OpenStudy (anonymous):

21 ft \[Area of a Trianlge = 1/2\times b \times h \]